OpenStudy (anonymous):

Find dp/dq using implicit differentiation: 2p^2*p*q=8p^2*q^2

OpenStudy (lgbasallote):

\[\large 2p^2pq = 8p^2q^2?\]

OpenStudy (anonymous):

yes that is right

OpenStudy (anonymous):

now how do i find the dp/dq of that equation using implicit differentiation?

OpenStudy (lgbasallote):

why dont you simplify first...

OpenStudy (anonymous):

would that be 4p*q?

OpenStudy (anonymous):

or p=4*q?

OpenStudy (turingtest):

second one

OpenStudy (anonymous):

so then what would the answer be?

OpenStudy (turingtest):

take the derivative with respect to q, then solve for dp/dq

OpenStudy (anonymous):

0?

OpenStudy (turingtest):

?

OpenStudy (anonymous):

would the answer be 0?

OpenStudy (turingtest):

no, the answer should be a function

OpenStudy (turingtest):

can you differentiate p=4q implicitly ?

OpenStudy (turingtest):

actually, in this case the answer is a number, but not 0

OpenStudy (anonymous):

4?

OpenStudy (turingtest):

yeah, I guess so boring problem :S

OpenStudy (anonymous):

its saying 4 isnt right:(

OpenStudy (callisto):

1/4?

OpenStudy (anonymous):

also not correct :(

OpenStudy (turingtest):

that's the only was my other guess, but I am starting to suspect a typo

OpenStudy (turingtest):

then there is a typo

OpenStudy (turingtest):

I was suspicious of the way the problem was printed...

OpenStudy (turingtest):

unless differentiating implicitly means we can't simplify, which is not a rule I knew about

OpenStudy (anonymous):

its 2p^2-pq=8p^2q^2

OpenStudy (anonymous):

that is the correct problem.. not the one i posted before danm it

OpenStudy (turingtest):

dude, you have a 8 where there is a - sign

OpenStudy (turingtest):

a * I mean...

OpenStudy (turingtest):

so back to the drawing board :P

OpenStudy (turingtest):

\[ 2p^2-pq=8p^2q^2\]

OpenStudy (anonymous):

yes thats right

OpenStudy (turingtest):

I'll let Callisto do it...

OpenStudy (anonymous):

thanks for tryin to help tt

OpenStudy (turingtest):

no prob :)

OpenStudy (callisto):

Find dp/dq using implicit differentiation: \[2p^2-pq=8p^2q^2\]\[p(2p-q)=8p^2q^2\]\[2p-q=8pq^2\] Diff. both sides wrt q \[\frac{d}{dq}2p-q=\frac{d}{dq}8pq^2\]\[2\frac{dp}{dq}-1=8q^2\frac{dp}{dq} + 16pq\]\[2\frac{dp}{dq}-8q^2\frac{dp}{dq}= 16pq+1\]\[(2-8q^2)\frac{dp}{dq}= 16pq+1\]\[\frac{dp}{dq}= \frac{16pq+1}{(2-8q^2)}\] It doesn't look right... Does it?

OpenStudy (anonymous):

noo it isnt right ;/

OpenStudy (callisto):

@TuringTest Sir~ Please help!!~ :|

OpenStudy (turingtest):

something is rotten in the state of Denmark... @Callisto I see no flaw in your reasoning, but perhaps there is a reason we can't simplify? I dunno, let me do it on paper...

OpenStudy (turingtest):

I get the same as @Callisto no matter how I slice it

OpenStudy (anonymous):

hmm maybe the problem is just a fluke who knows. i even tried an online calculator on it and it was saying that that answer it gave was also wrong while all the others i used it with came out correct thanks for your help anyways guys!

OpenStudy (callisto):

What was the 'correct' answer??

OpenStudy (callisto):

@drumjockey83

OpenStudy (anonymous):

never found it out

OpenStudy (turingtest):

\[2p^2-pq=8p^2q^2\]\[4pp'-p'q-p=16p^2q+16pq^2p'\]\[4pp'-p'q-16pq^2p'=16p^2q+p\]\[p'(4p-q-16pq^2)=16p^2q+p\]\[p'={16p^2q+p\over4p-q-16pq^2}\]

OpenStudy (turingtest):

so maybe when we simplified that cause to q become a 1 when when we took the derivative though it wasn't supposed too? I dunno, I wanna see why it makes a difference...

OpenStudy (turingtest):

I think that by simplifying we made it so we didn't have to use the product rule in the second term on the LHS, which caused us to drop the q I guess we have to consider that p could be zero, and therefor we can't divide by it???

OpenStudy (callisto):

Does that mean we can't simplify the given equation before differentiating it?

OpenStudy (turingtest):

Not a rule I was familiar with, but I guess we'll know if/when @drumjockey83 tells us if this is the right answer now

OpenStudy (anonymous):

correct turning test ! thank you so much!

OpenStudy (callisto):

The answer tells us all......